OpenStudy (anonymous):

simplify the square root of 8n-27 devided by 4-n

OpenStudy (anonymous):

kstatefan, do you know how to divide polynomials involving variables yet or no?

myininaya (myininaya):

-8-27/(4-n)

OpenStudy (anonymous):

I can show you if you need me to kstatefan. Just trying to assess how much I should go into detail

myininaya (myininaya):

what is 8n/(-n) simplify to

OpenStudy (anonymous):

-8 right?

myininaya (myininaya):

yes

myininaya (myininaya):

put that where the quotient goes and then multiply by the divisor to figure out what you need to subtract

OpenStudy (anonymous):

k

myininaya (myininaya):

oops i got the wrong remainder above

myininaya (myininaya):

-8+5/(4-n)

myininaya (myininaya):

so we have (8n-27)-(8n-32) which is ...

OpenStudy (anonymous):

this is exponential and logarithmic functions we arent doing this right

myininaya (myininaya):

oh thats the sqrt(8n-27)

myininaya (myininaya):

i didnt see the sqrt sorry

myininaya (myininaya):

so we aren't dividing. I would say it is already simplified

myininaya (myininaya):

how do you use log or exponentiation here

myininaya (myininaya):

exponential*

OpenStudy (anonymous):

and i said it wrong the funtion is actually\[\sqrt{(8^{n}-27)\div (4^{-n})}\]

OpenStudy (anonymous):

and i have no idea

myininaya (myininaya):

=\[\sqrt{8^n*4^n-27(4^n)}=\sqrt{32^n-27(4^n)}\]

myininaya (myininaya):

\[\sqrt{4^n}*\sqrt{(8^n-27})=2^n*\sqrt{(8^n-27)} \]